1. Consider a firm whose MRPL=60-0.25L. If the market wage is 20, how much labor does the firm use? (Answer format is number only, ie. 35, not L=35). (Answer is not 35).

2. The same firm as in Q9 is now offered a subsidy of $5 per worker by the government to increase employment. The market wage remains w=20. How much labor does the firm now hire?
Answer #1

1) Use MRPL = Wage rate as the demand and supply for labor. This implies 60 - 0.25L = 20. This gives L = (60 - 20)/0.25 = 160. Hence firm hires 160 labor units

2) Subsidy will increase demand for labor which means new MRPL is 60 - 0.25L + 5 or MRPL = 65 - 0.25L. This gives 65 - 0.25L = 20 or L = (65 - 20)/0.25 = 180. Hence firm hires 180 labor units after subsidy